OpenStudy (anonymous):

Glen is looking into buying a piece of land. He estimates that the land measures 100yd by850yd. The land is actually 140yd by 940yd. What is the percent error in the estimated area of the land? A.35.4 B.17.7 C.64.6 D. 32.3

OpenStudy (anonymous):

The formula for percent error is: (estimation − actualvalue)/(actual value) (times100percent)

OpenStudy (anonymous):

So the estimated area is: A = lw =100*850 =85 000 The actual area is: A = lw =140*940 =131 600 So the percent error in this estimate is: (85000 - 131600)/131600 (x100%) =about 35% So the answer is A

OpenStudy (anonymous):

I had that ,I wasnt sure. thank U!

OpenStudy (anonymous):

no problem